The perimeter of a rectangle is 420 inches. The width is 4 inches less than the length.
What is the length of the rectangle?

Answer by CharStar(110)

The formula for Perimeter = P 2(l+w)
Let Width = x-4
Let Length = x
420=2(x-4+x)
420=2x-8+2x
420=4x-8
Subract 8 from both sides
428=4x
107=x
Check your answer
P=2(l+w)
420=2(107-4+107)
420=2(210)
420=420
Hope this helps you